OpenStudy (dls):

A bodycovers 4m in 3seconds and 12m in 5 seconds.If motion is uniform..how far will it travel in nexxt 3 scs

OpenStudy (dls):

@yrelhan4

OpenStudy (dls):

sab kuch karlia..a=4 V=12m/s ab kya

OpenStudy (dls):

distance kaise nikalenge

OpenStudy (dls):

v^2-u^2=2as?

OpenStudy (yrelhan4):

haan shayad. answer nhi aa rha?

OpenStudy (dls):

usse 18m aega not sure i dont have the ans

OpenStudy (yrelhan4):

rest se hi start ho rha hai na?

OpenStudy (dls):

wahi to nahi pata?

OpenStudy (yrelhan4):

to fir V kyu likha hai. u likho. S=ut+1/2at^2 do cases me banake. u aur a nikalo. fir at the end of 8 seconds v nikalo, v=u+at se. fir next 3 seconds ke liye ye v=u', aur a to constant hi hoga. S=ut+1/2at^2 se distance nikal jayega.

OpenStudy (dls):

wahi to kiya..

OpenStudy (dls):

u=0 a ssume kiya bus

OpenStudy (dls):

wo chodo ye batao!! A train is running at a constant acc.The two ends of the train pass a telegraph pole with velocities v1 and v2.Find the velocity with which midpoint of the train passses the same pole.

OpenStudy (yrelhan4):

u=0 nhi krna hai.

OpenStudy (dls):

wtf

OpenStudy (yrelhan4):

yr ye last hai bas. mere itna saara kaam pada. fati padi hai meri. ruk.

OpenStudy (dls):

okay got it!

OpenStudy (yrelhan4):

ruk jao.

OpenStudy (yrelhan4):

dinner kar rha tha. dekh. v^-u^2=2as hota hai na. to v2^2-v1^2=2ax were x is length of train. let speed of middle point of train be v3 to v3^2-v1^2=(2ax/2) solve and find v3.

OpenStudy (dls):

kinda got it! thanks!

OpenStudy (yrelhan4):

solve krke dekhna. mereko is question ka answer bhi yaad hai. aa jayega.

OpenStudy (dls):

okay! abhi aage karr raha hoon :) hojaega ye...samaj agaya bus

OpenStudy (anonymous):

what !? motion is uniform.. then why is tehr acceleration!?

OpenStudy (dls):

nice question :P

OpenStudy (yrelhan4):

:O .. @DLS question me const acceleration likha hai kya? mere question me to yahi tha.

OpenStudy (dls):

1st ques or train one

OpenStudy (dls):

motion is uniform train is running at constant acceleration

OpenStudy (yrelhan4):

A train is running at a constant acc.The two ends of the train pass a telegraph pole with velocities v1 and v2.Find the velocity with which midpoint of the train passses the same pole. @Mashy bwahahaha

OpenStudy (anonymous):

what?! m talking about the first question!

OpenStudy (dls):

uniform motion

OpenStudy (anonymous):

how can it be uniform motion!???

OpenStudy (yrelhan4):

check the question dls.

OpenStudy (dls):

IDK,IDC

OpenStudy (anonymous):

you post crap questions :P

OpenStudy (dls):

given in NCERT

OpenStudy (anonymous):

uniform acceleration and i m pretty sure its the first 3 seconds.. then next 5 seconds.. and then asked us to find for the next 3 seconds .. so thats like super easy!

OpenStudy (dls):

yeah...

OpenStudy (anonymous):

so you form two equations and solve them!

OpenStudy (dls):

arre hogaya na :/ badbadai ja raha hai :P

OpenStudy (anonymous):

ok ok good :)

OpenStudy (anonymous):

train problem also done?!..

OpenStudy (dls):

yes!

OpenStudy (anonymous):

now i saw the question is closed.. lolz!

OpenStudy (dls):

.___.

OpenStudy (yrelhan4):

lolz? laugh out loud zebras? @Mashy

OpenStudy (dls):

lol